OpenStudy (kingethan8500):

The volume of a cone is 392.5 cubic inches. The height of the cone is 15 inches. What is the radius of the cone, rounded to the nearest inch? (Use π = 3.14.) 3 5 13 25

OpenStudy (astrophysics):

\[V = \frac{ 1 }{ 3 }\pi r^2\] solve for and plug in the values, V=392.5 in^3, h = 15in, pi = 3.14

OpenStudy (astrophysics):

solve for r*

OpenStudy (kingethan8500):

how

OpenStudy (astrophysics):

Lets first solve for r^2 so multiply both sides by 3 and then divide by pi

OpenStudy (kingethan8500):

sorry im lost i dont understand

OpenStudy (astrophysics):

\[\color{red}{3}V = \frac{ 1 }{ 3 } \pi r^2 *\color{red}{3}\] notice the 3's on the left cancel each other out making a 1, so we have \[3V= \pi r^2\] then divide both sides by pi \[\frac{ 3V }{ \color{red}{\pi} } = \frac{ \pi r^2 }{ \color{red}{\pi} }\] again on the right sides pi's cancel each other out giving a 1, \[r^2 = \frac{ 3V }{ \pi }\] now we have that squared we want to get rid of to do that we can square root both sides because \[\sqrt{x^2} = x\]

OpenStudy (retireed):

I think we know the answer, but where would the poster plug in the 15 inch height in your formula?

OpenStudy (kingethan8500):

so the answer is 3?